a triangle ABC,if  r1 is the exradius of the excircle opposite to A,r2 is the exradius of the excircle opposite to B and r3is the exradius of the excircle opposite to C....


In prove that (ar1+br2+cr3)/abc=(1/r)-(1/2R)


where r is the inradius and R is the circum radius.solytion of trainle


Grade:12

2 Answers

Rahul Kumar
12 years ago

Hi Anil,

You font size is very small, please increase it and post again so that we can reply that

Regards

anil kumar jha
12 years ago

properties of triangles. in the following problem, In a triangle ABC,if r1 is the exradius of the excircle opposite to A,r2 is the exradius of the excircle opposite to B and r3is the exradius of the excircle opposite to C.... prove that (ar1+br2+cr3)/abc=(1/r)-(1/2R) where r is the inradius and R is the circum radius. i tried using these formulae: r1=(area of triangle)/(s-a) r1=stanA/2 but i cudnt solve it.how should i start?